# Honeybee: Decentralized Peer Sampling with Verifiable Random Walks for Blockchain Data Sharding

### v1

@@ 4.1. Algorithm Overview
Honeybee  is a fully decentralized p2p algorithm that tackles the data availability sampling problem on the network level by conducting secure near-uniform [^4] Note that near-uniform is well-defined for peer sampling. sampling of peers from the entire network. Honeybee  is inspired by random walks on finite groups and mixing Markov chains (Aldous, 1983; Lovász, 1993). A Honeybee  node maintains three tables: an address table, an encounter table, and a connection table. In the address table, a Honeybee  node stores peer data for peers with whom the node has peering agreements. The address table consists of two sub-tables: an outgoing address table that contains data for at most $n_{out}$ peers and an incoming address table that contains data for at most $n_{in}$ peers. If a node $A$ sends a request to peer with node $B$, and node $B$ accepts the request, node $A$ will add node $B$ to $A$’s outgoing address table, while node $B$ will add node $A$ to $B$’s incoming address table. The peering agreement in Honeybee  is bilateral - node $A$ having node $B$ in $A$’s address table would suggest that node $B$ having node $A$ in $B$’s address table. To be specific, when a node $A$ has node $B$ in its address table, node $A$ stores the following data in its address table: $B$’s node ID, $B$’s address table snapshot, $B$’s IP address, $B$’s port information, and a peering agreement signed by both $A$ and $B$ with a timestamp. A node’s address table snapshot contains everything in its address table except for its peers’ address table snapshots. A node’s encounter table is a finite FIFO list of size $n_{e}$ that contains the address table snapshots of a random subset of the nodes it has interacted with via random walks. A node’s connection table contains the peers with whom the node communicate.
(a) Overview of Honeybee VRW
(b) Illustration of Honeybee VRW
(c) Illustration of Honeybee TCC
Figure 1. Honeybee illustrations: (a) shows an example of a Honeybee  verifiable random walk sampling, and node $C$ replaces node $B$ in node $A$’s outgoing address table; (b) shows how a verifiable random walk (VRW) works; (c) shows how a table consistency check (TCC) works.
To join the Honeybee  network, a node first add some of the bootstrap nodes for an initial configuration. The bootstrap nodes are a small group of trustworthy nodes that not only participate in the network but also assist other nodes with configurations. In Honeybee, all nodes follow the blockchain for the shared pseudo-randomness (i.e., block header), and the shared pseudo-randomness determines the nodes that are eligible to perform sampling at the current moment (currently eligible nodes). To sample peers from the network in a near-uniform way, a currently eligible node conducts a random walk with the path determined by the shared pseudo-randomness. A random walk message carries its initiator node information, destination node information, and the random walk path information. During a random walk, the node on each hop of the random walk conduct table consistency checks with the initiator node. When a random walk message reaches its destination node, the message serves as a request for the random walk initiator node to peer with the destination node. If the destination node accepts the request, the initiator node and the destination node add each other to their address tables and establish a peering agreement. For a node, we call the time between successive roles as a currently eligible node as an epoch (round). In Algorithm 1, we present the overall template for outgoing address table update of Honeybee. We give the definitions of verifiable random walks and table consistency checks in §4.2 and §4.3. In Fig. 1a, we give an example of a Honeybee  verifiable random walk sampling. Fig. 1b shows how a verifiable random walk works. Fig. 1c shows how a table consistency check works.
@@
input : peers $\Gamma^{v}_{\text{curr}}$ in outgoing address table of current epoch; pseudo-randomness seed $\mathcal{R}$ indicating whether node $v$ is a currently eligible node; $v$’s secret key $SK_{v}$; epoch counter $\varrho_{\text{curr}}$ enumerating the number of epochs since $v$ joins the network; outgoing address table size $n_{out}$; system-defined address table inconsistency threshold $\tau$; bootstrap node addresses $b$;
output : updated set of peers $\Gamma^{v}_{\text{next}}$ for next epoch; updated epoch counter $\varrho_{\text{next}}$ for next epoch;
/* Perform a random walk if $v$ is a currently eligible node */
@@
       $p\leftarrow\textsc{GetPathLength}(\mathcal{R},v,SK_{v})$ $p_{c}\leftarrow 0$ /* Hop counter of the random walk */
       $u^{*}\leftarrow v$ /* Node on current hop of the random walk */
       $d\leftarrow\emptyset$ /* $d$ stores the node on the first hop of the random walk */
@@
             /* VRF() returns a pseudo-random number $i$ and a proof $\pi_{v}$/
             $(i,\pi_{v})\leftarrow VRF(\mathcal{R},\varrho_{\text{curr}},p_{c},u^{*},SK_{v})$ /* $\Gamma^{u^{*}}_{\mathrm{curr}}(i)$ returns i-th node in $\Gamma^{u^{*}}_{\mathrm{curr}}$ and its address table snapshot */
             $(u^{*},\omega)\leftarrow\Gamma^{u^{*}}_{\mathrm{curr}}(i)$ $p_{c}\leftarrow p_{c}+1$ if $p_{c}=1$ then
                   $d\leftarrow u^{*}$
@@
            $\omega^{*}\leftarrow\textsc{RequestAddrTable}(u^{*})$ if $\textsc{Diff}(\omega,\omega^{*})>\tau$ then
                   $\textsc{IssueFraudProof}(u^{*},\omega,\omega^{*})$
             end if
       end while
      /* GetDestResponse() returns the response $\varphi$ of the destination node */
       $\varphi\leftarrow\textsc{GetDestResponse}(u^{*})$ if $\varphi$ indicates $u^{*}$ accepts request from $v$ then
             $\Gamma^{v}_{\text{next}}\leftarrow\Gamma^{v}_{\text{curr}}\backslash\{d\}\cup\{u^{*}\}$
       end if
      /* Add new peers if outgoing address table is not full */
       if Size$(\Gamma^{v}_{\text{curr}})$ $<$ $\mathrm{n_{out}}$  then
             $\Gamma^{v}_{\text{next}}\leftarrow\Gamma^{v}_{\text{curr}}\cup\textsc{RequestNewPeers}(b,n_{out}-\textsc{Size}(\Gamma^{v}_{\text{curr}}))$
       end if
      $\varrho_{\text{next}}\leftarrow\varrho_{\text{curr}}+1$
end if
@@
Algorithm 1 Algorithm outline for updating entries of outgoing address table of node $v$ in each epoch.
4.2. Verifiable Random Walks
A Honeybee  node performs a random walk to replace one of its outgoing table peers when it learns that it becomes a currently eligible node from the pseudo-randomness seed (i.e., block header). However, a random walk is susceptible to adversarial routing. If a random walk from a honest node meets a dishonest node, the dishonest node can route the random walk to another dishonest node and eventually guide the random walk to a dishonest destination. To ensure our random walks are truly random, we employ a verification mechanism in Honeybee’s random walks (VRW). When a node $A$ learns that it is a currently eligible node, it derives a verifiable pseudo-random number and a proof $\pi_{A}$ with its secret key using the following inputs: the current block header, its epoch count, its random walk’s current hop count, and the public key of the node on its current hop. The generated verifiable pseudo-random number informs node $A$ where to proceed with the next hop of its random walk. Along with the signature of the node on the current hop, the nodes on the next hop of the random walk can verify the validity of the random walk. The procedure is repeated for each hop until the end of the random walk. At node $B$, node $A$ requests the address table snapshot of the next hop - node $C$. When node $A$ proceeds to node $C$, it requests the current address table of node $C$ and compare it with the snapshot $A$ receives from $B$. If the difference is higher than a threshold $\tau$, $A$ may issue a fraud proof against $C$. The threshold $\tau$ is predefined by the system. Since nodes are required to update their address table snapshots with their neighbors promptly, node $C$ cannot refute the VRW fraud proof.
4.3. Table Consistency Checks
@@
A dishonest node can store multiple copies of address tables, and use different address tables to handle different requests from different nodes for various purposes (e.g., traffic attraction, adversarial routing, etc.). We refer to the situation where a node uses more than one address table as ”equivocal tables.” To prevent equivocal tables, we employ table consistency checks (TCC). A node $A$’s random walk message carries $A$’s address table and encounter table. For each hop of the random walk, the node $B$ at that hop can compare its address table and encounter table with those of node $A$. If there is an overlapping node $C$ in $A$’s tables and $B$’s tables, $A$ and $B$ will further examine the address table snapshots of node $C$ to check if there is any difference. If the difference is higher than a threshold $\tau_{t}$, $A$ and $B$ will issue a fraud proof against node $C$ and the system may choose to slash node $C$. Similar to VRW, the threshold $\tau_{t}$ is predefined by the system. If the ratio of different peers in two address table snapshots of node $C$ within time $t$ exceeds $\tau_{t}$, then we consider the evidence is significant enough for nodes to issue a fraud proof for node $C$. Node $C$ can refute the TCC fraud proof against it only if it can provide legitimate random walk history that can explain the difference.
@@ 5.1. Experimental Setup
We describe the baseline algorithms in §5.1.1, network layouts in §5.1.2, adversary configurations §5.1.3, and adversary strategies in §5.1.4.
@@ 5.1.1. Baselines
Since the main goal of Honeybee  is to conduct near-uniform peer sampling in p2p networks under the data availability setting, we compare Honeybee  with two arguably most potential candidate algorithms under such setting (Neu, 2022) - Kademlia (Maymounkov and Mazieres, 2002) and GossipSub (Vyzovitis et al., 2020).
  • Kademlia: Kademlia is one of the most popular p2p protocol in today’s Internet. It is used in various systems including Ethereum, Swarm, Storj, IPFS, etc. A Kademlia node has a binary node ID that is randomly assigned when the node joins the network. To route messages, each Kademlia node has a routing table consists of ”k-buckets”, and the number of k-buckets it has is equal to the length of its node ID. The $i$-th k-bucket of node A stores peers with node IDs that share the first $i-1$ bits with node A. A Kademlia node discovers new peers mainly with lookups. When node $A$ performs a lookup on node ID $B$, $A$ sends the lookup message to the neighbor(s) whose node ID is closest to $B$ in terms of XOR distance. The neighbor(s) and the nodes on the subsequent hops repeat this procedure until they find the closest node(s) to $B$. For more details about Kademlia, we refer the reader to the Kademlia paper (Maymounkov and Mazieres, 2002).
@@ 5.1.2. Network
We built a discrete-event network simulator using Python based on the model description in §3.1. We simulate three p2p networks - the Kademlia network, the GossipSub network, and the Honeybee  network. Each of these networks consists of 16,384 ($=2^{14}$) nodes. The Kademlia network simulates the vanilla Kademlia network as discussed in §5.1.1. The Kademlia nodes each has 14 k-buckets and each k-bucket has size of 3 (note that some buckets may never have enough peers to reach the size limit), and we use $\alpha=3$ for all lookups. The GossipSub network simulates one overlay of the vanilla GossipSub network as discussed in §5.1.1. We simulate peer discovery in GossipSub with peer exchanges. We also simplified the scoring function by giving every peer the same score to fit GossipSub into the data availability sampling background. The scoring function from GossipSub may not help honest nodes in the data availability sampling setting. The scoring function uses parameters such as time in mesh, first message deliveries, and mesh message delivery rates. With these parameters, dishonest nodes can strategically exploit the scoring function and behave very well in terms of scores before eclipsing an honest node. The GossipSub nodes use $D\_high=12$, $D=8$, and $D\_low=6$, which are the same as the values used in the vanilla GossipSub paper. The Honeybee  network simulates Honeybee  as discussed in §4. For the Honeybee  network, 17 nodes ($\approx 1\%$) are the bootstrap nodes. Similarly, the Honeybee  nodes each uses an address table size of 24. The address table of a Honeybee  node consists of the outgoing address table and the incoming address table, each contains at most 12 nodes. In all three networks, the routing table of a node is defined as all the peers whom it knows and has access to (i.e., stores their IP addresses), and we assume there is no churn from the joining and leaving of nodes.

### v4

@@ 3.1. Algorithm Overview
We propose Honeybee. Honeybee is a fully decentralized p2p algorithm that tackles the node sampling problem by conducting secure near-uniform sampling of peers from the entire network. [^3] Note that near-uniform is defined for peer sampling in §1 and §2.2. In Honeybee, a node samples an address by carrying out a random walk of sufficient length on the network. To ensure the samples are uniformly random, the graph on which the random walk is conducted is designed to be regular resulting in a doubly-stochastic probability transition matrix with a uniform stationary distribution for the Markov chain (Aldous, 1983; Lovász, 1993). The graph on which the random walks are executed is different from the network’s communication overlay, which can have an arbitrary topology per the application’s specifications. To ensure that the regularity of the random walk graph is maintained and random walks are carried out correctly, Honeybee relies on two key ideas: (1) random walk paths (and their lengths) are decided pseudo-randomly using verifiable random functions and public randomness source (details in §3.3); (2) addresses sampled by a node are independently and secretly cross-verified by other nodes. If an inconsistency is discovered, a fraud proof is issued to the network (details in §3.4).
3.2. Address Table Design
A Honeybee node $v$ maintains three tables: an address table, an encounter table, and a connection table. The address table holds the addresses of the nodes most recently sampled by $v$. Entries in the address table are regularly replaced by fresh samples through Honeybee’s sampling algorithm. The encounter table holds addresses of nodes most recently visited by $v$ as part of Honeybee’s random walks. Addresses in the encounter table help to bootstrap random walks when the address table does not contain sufficient addresses. A node’s connection table contains peers with whom the node communicates. The connection table can be a subset of addresses from the current or past address table.
When a node joins the Honeybee network, it first contacts a bootstrapping server to receive an initial set of random peer addresses for its address table. Bootstrapping servers are a small group of trustworthy nodes that assist other nodes with their initial configurations. We assume the bootstrapping server has the capability to deliver uniform peer samples to requesting nodes. Such a capability can be achieved by letting the bootstrapping nodes also perform periodic peer sampling, or by other means.
The address table for node $v$ consists of two sub-tables: an outgoing address table $\Gamma^{v}_{out}$ that contains data for at most $n_{out}=k/2$ peers and an incoming address table $\Gamma^{v}_{in}$ that contains data for at most $n_{in}=k/2$ peers. The outgoing address table contains node addresses that $v$ has sampled, while the incoming address table contains the addresses of nodes which have sampled $v$. In Honeybee, for two nodes $u,u^{\prime}$, if $u^{\prime}\in\Gamma^{u}_{out}$ then $u\in\Gamma^{u^{\prime}}_{in}$ and vice-versa. To ensure this property, if a node $u$ samples a node $u^{\prime}$, it first sends a request to $u^{\prime}$. If node $u^{\prime}$ accepts the request, $u$ adds $u^{\prime}$ to the outgoing address table of $u$, while $u^{\prime}$ adds $u$ to the incoming address table of $u^{\prime}$. This process involves $u$ and $u^{\prime}$ signing a ‘peering agreement’ with an expiration time (e.g., $k\Delta$). If the peering agreement is not renewed before its expiry, then $u$’s address in the address table of $u^{\prime}$ (and vice-versa) is considered invalid. Address table updates in Honeybee are thus bilateral. Conversely, when $u$ updates its address table to replace $u^{\prime}$ by another address, $u$ communicates the update to $u^{\prime}$ so that $u^{\prime}$ can remove $u$ from its address table. Addresses that are not removed cooperatively from a node’s address table naturally time out and become verifiably invalid.
If $u$ has $u^{\prime}$ in its address table, $u$ stores the following data for $u^{\prime}$ in the address table: public key of $u^{\prime}$ (certified by a bootstrapping server, or by a decentralized certificate authority), a snapshot of the address table of $u^{\prime}$, IP and port address of $u^{\prime}$, and a peering agreement signed by both $u$ and $u^{\prime}$ with an expiry timestamp. A node’s address table snapshot contains everything in its address table except for its peers’ address table snapshots. These snapshots help to verify the correctness of random walks as discussed next. [^4] We refer to the shared copy of an address table as a snapshot because, while it may be susceptible to modification or tampering, it can also serve as evidence against the table’s owner later, as it is digitally signed.
@@ 3.2. Address Table Design
input : peers $\Gamma^{v}_{\text{curr}}$ in outgoing address table of current epoch; shared pseudo-randomness seed $\mathcal{R}$; $v$’s secret key $SK_{v}$
output : updated set of peers $\Gamma^{v}_{\text{next}}$ for next epoch
@@
       /* $v$ infers the random walk path length $p$ from $\mathcal{R},v,SK_{v}$. */
       $p_{c}\leftarrow 0$ /* Hop counter of the random walk */
       $u^{*}\leftarrow v$ /* $u^{*}$ stores node on current hop of the random walk */
       /* $d$ stores the index of node on first hop of the random walk */
@@
             /* $v$ performs one hop of the iterative verifiable random walk and updates $p_{c},u^{*}$; $v$ issues a fraud proof against $u^{*}$ when snapshots from VRW or TCC for the same node $u^{*}$ differs materially (more details in Appendix A) */
       end while
      $\Gamma^{v}_{\text{next}}\leftarrow\Gamma^{v}_{\text{curr}}\backslash\{\Gamma^{v}_{\text{curr}}(d)\}\cup\{u^{*}\}$
@@
Algorithm 1 Algorithm outline for updating entries of the outgoing address table of node $v$ in each epoch.
In Algorithm 1, we present the overall template for outgoing address table update of Honeybee.
3.3. Verifiable Random Walks (VRW)
In each round of Honeybee, a fraction $\eta$ of nodes in the network attempt to obtain a fresh sample for their outgoing address table. Restricting address table updates to only a fraction of the nodes avoids excessive churn in the address tables which in turn helps maintain the stationary distribution of the random walks to be uniform. A node determines its eligibility to conduct a random walk during a round by computing a verifiable random function (VRF) over the shared public randomness for the round. When a node $u$ becomes eligible during a round, it performs a random walk to replace one of its outgoing table peers.
A key challenge in performing a random walk is adversarial routing: if a random walk from an honest node meets a dishonest node, the dishonest node can route the random walk to another dishonest node and eventually guide the random walk to a dishonest destination. To avoid adversarial routing, in Honeybee the length $l$ and path $p$ of a node $u$’s random walk are determined pseudo-randomly using the round’s public randomness and the VRF. The input to the VRF includes the current round number and public randomness value. The path $p$ of the random walk is implicitly specified as a sequence $x_{1},x_{2},\ldots,x_{l}$ of indices where each $x_{i}\in\{1,2,\ldots,k\}$ for all $i$. To conduct a random walk, the node $u$ visits each node of $p$ sequentially. Supposing node $u$ visits node $u^{\prime}$ for hop $i-1$, and moves to node $u^{\prime\prime}$ for hop $i$. Recall that $u^{\prime}$ has a snapshot of the address table of $u^{\prime\prime}$, and vice-versa. When $u$ contacts $u^{\prime\prime}$, node $u^{\prime\prime}$ first verifies that the most recent hop was executed correctly, i.e., the $x_{i-1}$-th entry in the address table of $u^{\prime}$ is $u^{\prime\prime}$. Node $u^{\prime\prime}$ also verifies whether the current hop number for $u$ is not greater than $l$. To do this, $u$ submits snapshots of the address tables from each of the $i-1$ nodes visited so far in the walk to $u^{\prime\prime}$. The address table snapshots collected by $u$ are signed by their respective owners. A table consistency check algorithm (§3.4) verifies that the signed address table snapshots are valid, and thus verifies the random walk length. The walker $u$ then requests node $u^{\prime\prime}$ to provide the IP address of the $x_{i}$-th address within the address table of $u^{\prime\prime}$. To verify whether the received address is correct, $u$ checks the address received from $u^{\prime\prime}$ against the snapshot of the address table of $u^{\prime\prime}$ received from $u^{\prime}$. If the address is verified to be correct, node $u$ then proceeds to the $i+1$-th step of the walk. Finally, when $u$ reaches the last node of the walk it requests the node to be added as a new sample within $u$’s table. If the requested node accepts the request, the initiator node $u$ and the destination node add each other to their address tables and establish a peering agreement. A snapshot of $u$’s updated address table is shared with all peers in its address table. For a node, we call the time between successive roles as an eligible node as an epoch.
(a) Honeybee Random Walk
(b) Honeybee VRW
(c) Honeybee TCC
Figure 1. Honeybee illustrations: (a) an example of Honeybee verifiable random walk (VRW) sampling with three hops; (b) how VRW works from hop to hop; (c) how table consistency check (TCC) works among nodes.
In Fig. 1a, we give an example of Honeybee verifiable random walk sampling with three hops, and the destination node replaces the node on the first hop in the outgoing address table of the initiator node at the end of the sampling. Fig. 1b shows how a verifiable random walk works.
3.4. Table Consistency Checks (TCC)
@@
A dishonest node can store multiple address tables, and use different address tables to handle different requests from different nodes for various purposes (e.g., traffic attraction, adversarial routing, etc.). We refer to the situation where a node uses more than one address table as “equivocal tables.” To prevent equivocal tables, we employ table consistency checks (TCC) in Honeybee. When a node $u$ visits a node $u^{\prime}$ during a random walk, $u$ and $u^{\prime}$ determine whether there are any addresses common to the address and encounter tables of $u$ and $u^{\prime}$. If an overlapping node $u^{\prime\prime}$ is found in the tables of both $u$ and $u^{\prime}$, subsequently $u$ and $u^{\prime}$ will further compare the address table snapshots of node $u^{\prime\prime}$ as stored by $u$ and $u^{\prime}$ to check if there is any difference. If the difference is higher than a threshold $\tau_{t}$ (i.e., more than $\tau_{t}$ addresses are different in the snapshot versions of $u$ and $u^{\prime}$), then $u$ and $u^{\prime}$ will issue a fraud proof against node $u^{\prime\prime}$ and the system may choose to slash node $u^{\prime\prime}$. $\tau_{t}$ is predefined by the system, and $\tau_{t}$ increases as $t$ (time difference between two snapshot versions) increases. $t$ is bounded from above by a system-defined parameter $t_{max}$ such that nodes only need to store a limited history of their past random walks. [^5] Raising $t_{max}$ improves system security but increases storage demands on nodes. In the analysis (Appendix B), we use $t_{max}=\frac{k}{2\eta}$. If the ratio of different peers in two address table snapshots of node $u^{\prime\prime}$ within time $t$ exceeds $\tau_{t}$, then we consider the evidence is significant enough for nodes to issue a fraud proof for node $u^{\prime\prime}$. Node $u^{\prime\prime}$ can refute the TCC fraud proof against it only if it can provide legitimate random walk history that can explain the difference. Fig. 1c gives an example of a table consistency check.
@@ 4. Evaluation
To understand the effectiveness of Honeybee theoretically, we conduct an analysis of Honeybee and include the results in Appendix B.
@@ 4.1. Experimental Setup
Here, we describe our experimental setup, which includes the baseline algorithms, network layouts, adversary configurations , and adversary strategies.
@@ 4.1.1. Baselines
Since the main goal of Honeybee is to conduct near-uniform peer sampling in p2p networks, we compare Honeybee with two arguably most potential candidate algorithms under such setting (Neu, 2022) - Kademlia (Maymounkov and Mazieres, 2002) and GossipSub (Vyzovitis et al., 2020). We briefly address them below.
  • Kademlia: Kademlia is one of the most popular p2p protocol in today’s Internet. It is used in various systems including Ethereum, Swarm, Storj, IPFS, etc. A Kademlia node has a binary node ID that is randomly assigned when the node joins the network. To route messages, each Kademlia node has a routing table consists of sub-tables called “k-buckets,” and the number of k-buckets it has is equal to the length of its node ID. The $i$-th k-bucket of node A stores peers with node IDs that share the first $i-1$ bits with node A. A Kademlia node discovers new peers mainly with lookups. When node $A$ performs a lookup on node ID $B$, $A$ sends the lookup message to the neighbor(s) whose node ID is closest to $B$ in terms of XOR distance. The neighbor(s) and the nodes on the subsequent hops repeat this procedure until they find the closest node(s) to $B$. For more details about Kademlia, we refer the reader to the Kademlia paper (Maymounkov and Mazieres, 2002).
    In our evaluation, the Kademlia nodes each has 14 k-buckets and each k-bucket has size of 3 (note that some buckets may never have enough peers to reach the size limit) such that the baseline algorithms have similar memory space sizes, and we use $\alpha=3$ for all lookups to make Kademlia more efficient (i.e., we send out three parallel queries to perform efficient lookups).
@@ 4.1.2. Network
We built a discrete-event network simulator using Python based on the model description in §2.1. We simulate three p2p networks - the Kademlia network, the GossipSub network, and the Honeybee network. Each of these networks consists of 16,384 ($=2^{14}$) nodes, and 17 of them ($\approx 0.1\%$) are the bootstrap nodes. The Kademlia network simulates the Kademlia network as discussed in §4.1.1, and the GossipSub network simulates one overlay of the GossipSub network as discussed in §4.1.1. The Honeybee network simulates Honeybee as discussed in §3. Similar to the baselines, the Honeybee nodes each uses an address table size of 24. The address table of a Honeybee node consists of the outgoing address table and the incoming address table, each contains at most 12 nodes. In all three networks, a node’s routing table is defined as the peers whom it knows and has access to (i.e., stores IP addresses), and we assume there is no churn from the joining and leaving of nodes.